OpenStudy (christos):

cscθ = 1/cotθ is this formula correct or not?

OpenStudy (anonymous):

\[\cscθ = \frac{ 1 }{ \cotθ }\] \[\csc(\theta) = \tan(\theta)\]

OpenStudy (anonymous):

what is the formula for exactly?

OpenStudy (anonymous):

@Christos

OpenStudy (christos):

its for a problem, dude csc(Θ) cant be equal to tan(Θ) !!!

OpenStudy (anonymous):

u sure?